OpenStudy (anonymous):

125-y^3 -what is the answer ? from binomial to trinomial i need help right now

OpenStudy (anonymous):

125=5^3 y^3 use formula (a^3-b^3) here a=5 and b=y

OpenStudy (anonymous):

\[a^3 - b^3 = (a-b)(a^2 + b^2 + ab)\]

OpenStudy (anonymous):

Here a = 5 and b = y Put these values in the formula...